Sum to n terms of the series
1.3.5.7 + 3.5.7.9+............

Grade:11

8 Answers

noogler
8 years ago
first find the nth term of the series
1,3,5,7,........are in first place …........nth term=1+(n-1)2=2n-1                    nth term=a+(n-1)d
3,5,7,9.........are in second place.........nth term=3+(n-1)2=2n+1
similarly 3rd and 4th places nth terms are 2n+3,2n+5 respectively
now the entire nth term is........(2n-1)(2n+1)(2n+3)(2n+5)
exp it and apply summation..................................................................................  
raju
8 years ago
∑(2n-1)(2n+1)(2n+3)(2n+5)
charan
8 years ago
 
 
∑(2n-1)(2n+1)(2n+3)(2n+5)
OUTPASS
8 years ago
∑(2n-1)(2n+1)(2n+3)(2n+5)
HARINATH
8 years ago
∑(2n-1)(2n+1)(2n+3)(2n+5)
salvathshaik
8 years ago
∑(2n-1)(2n+1)(2n+3)(2n+5
saibhargav
8 years ago
∑(2n-1)(2n+1)(2n+3)(2n+5)
Raunak JhunJhunWala
6 years ago
Well it can be done in more easier process by Vn method. nth term=2n-1*2n+1*2n+3*2n+5. Then apply summation properties to get answer....